# WoVoGen: World Volume-aware Diffusion for Controllable Multi-camera Driving Scene Generation

### v1

@@ B.1 Downstream tasks
original 34.9 69.2 20.4 17.9 124.6 28.6
+ generated 36.2 68.6 20.1 15.7 123.4 28.1
Table 2: The enhancement brought about by our generated data in 3D object detection is tested using the BEVDet model [13]. The validation is on the vehicle classes of cars, trucks, and buses.
@@
B.2 Ablation studies
Effectiveness of object guidance We implement object guidance to emphasize the location of the generated target. The qualitative comparison in Figure 8 illustrates the ablation study, while the quantitative comparison, using the metric that more accurately depict the alignment among image instances, i.e., CLIP FID [17], is presented in Table 3.
Method $FID_{clip}\downarrow$
w/o object guidance 20.3
w/ object guidance 12.8
Table 3: Quantitative comparison of image generation quality on the NuScenes validation set with and without object guidance
Clearly, the integration of object guidance results in a more precise alignment of the object within the generated image with the target image. This refinement corresponds to a notable reduction of 7.5 in the metric, $FID_{clip}$.
Effectiveness of temporal finetuning To verify the effectiveness of the video consistency module, we illustrate the simulation outcomes of the front and back cameras in Figure 9 after removing the temporal finetuning.
@@ B.2 Ablation studies
B.3 Additional single-frame image generation results
@@
Figure 8: Qualitative ablation study on object guidance. The incorporation of object guidance significantly improves the alignment of the generated image’s object with the target image.
@@
B.4 Additional video generation results
More generated videos along with the corresponding generated world volume are shown in Figure 11, Figure 12, Figure 13 and Figure 14.
The world model accurately predicts future world evolution and generates a comprehensive world volume. Leveraging world volume-aware diffusion, it produces high-quality videos that maintain cross-frame consistency and coherence across multiple cameras.

### v4

@@ 0.A.3 Downstream tasks
original 17.45 34.9 69.2 20.4 17.9 124.6 28.6
+ generated 18.10 36.2 68.6 20.1 15.7 123.4 28.1
Table 3: The enhancement brought by our generated data for 3D object detection. The evaluation is on the vehicle classes of cars, trucks, and buses.
@@
0.A.4 Ablation studies
Effectiveness of temporal finetuning
To verify the effectiveness of the video consistency module, we illustrate the simulation outcomes of the front and back cameras in Figure 9 after removing the temporal finetuning.
@@ 0.A.4 Ablation studies
0.A.5 Video results
We show the video results in file 10410_video.mp4.
0.A.6 Additional single-frame image generation results